OpenStudy (anonymous):

find the equation of the line having the given slope and containing the given the pair of point (-7,-2) and (-4,-3)

OpenStudy (anonymous):

1) find the slope: m=(y1-y)/(x1-x)

OpenStudy (anonymous):

1) m=(-2-(-3))/(-7-(-4))=1/(-3)=-1/3

OpenStudy (anonymous):

now use any given point to write your equation: y-(-2)=-1/3 *(x-(-7)) y+2=-x/3 -7/3 y=-x/3 -7/3-2 y=-1/3 x -13/3